OpenStudy (anonymous):

Compute the area of a triangle with sides. 10 x 10 x 16 a = ___ square units

OpenStudy (anonymous):

axbxc

OpenStudy (anonymous):

1,600

hartnn (hartnn):

nopes, use Heron's formula

OpenStudy (anonymous):



OpenStudy (anonymous):

10 10 8

OpenStudy (anonymous):

I like to picture it then you can figure it out often without remembering a formula.

OpenStudy (anonymous):

10^2x10^2x8^2

hartnn (hartnn):

i have a doubt @just.chris how do u know that the perpendicular line drawn also bisects the side?? how do u get 8 ? sorry if i am interrupting.....

OpenStudy (jiteshmeghwal9):

\[\LARGE{\color{Red}{Heron's \space Formula:-}}\]\[\LARGE{\color{green}{\sqrt{s(s-a)(s-b)(s-c)}}}\]where,\(\Large{\color{blue}{s=\frac{1}{2}(a+b+c)}}\)

OpenStudy (jiteshmeghwal9):

find s first of all :)

OpenStudy (anonymous):

10x10x16

OpenStudy (jiteshmeghwal9):

& then put the values in the formula to get ur answer :)

OpenStudy (anonymous):

10 x 10 x 16 is an equilateral triangle.

hartnn (hartnn):

@just.chris no, its not equilateral.......

OpenStudy (jiteshmeghwal9):

isosceles triangle

OpenStudy (jiteshmeghwal9):

becoz its two sides are equal dude ;)

OpenStudy (anonymous):

it has 2 sides the same

OpenStudy (jiteshmeghwal9):

& one side is diffrent

OpenStudy (anonymous):

dividing it through the mid point of the 16 side to the angle between the 10 sides gives you two right triangles.

OpenStudy (anonymous):

so im multipling 10x10x16 right

OpenStudy (anonymous):

Oops sorry isosceles. But you still get the two right triangles.

OpenStudy (jiteshmeghwal9):

why ??

OpenStudy (jiteshmeghwal9):

which formula r using ??

OpenStudy (anonymous):

trying to find s

OpenStudy (jiteshmeghwal9):

here, a=10 b=10 c=16\[s=\frac{10+10+16}{2}\]\[s={{36 \over 2}=18}\]

OpenStudy (jiteshmeghwal9):

\[\LARGE{\sqrt{18(18-10)(18-10)(18-16)}}\]now solve :)

hartnn (hartnn):

@just.chris if can please elaborate on how its right triangle, it would be great, because you would have found much simpler way to find the area.

OpenStudy (anonymous):

jiteshmeghwal9 we should do this not in Claire4Christ question unless she is interested in the discussion.

OpenStudy (anonymous):

im fine

OpenStudy (anonymous):

I don't mind it lol

OpenStudy (jiteshmeghwal9):

LOL if he/she is not interested then why did she ask the question ??

OpenStudy (jiteshmeghwal9):

@just.chris

OpenStudy (anonymous):

324-180+324-180+324-288

OpenStudy (anonymous):

Angle bisector theorem. But it makes sense too looking at the attached diagram.

OpenStudy (anonymous):

I always try to get everything to right triangles, maybe it is a weakness of mine, but if I can I can do almost everything just by remembering a^2 + b^2 = c^2.

OpenStudy (anonymous):

20+20+256

OpenStudy (perl):

hey hartn, in the last problem we did with teh pulley, so basically dh/dt is the same as dy/dt

OpenStudy (perl):

, yes one moment

OpenStudy (perl):

we have a nice formula for isosceles triangle, area , we can use pythagorean theorem

OpenStudy (perl):

here is a problem very similiar http://answers.yahoo.com/question/index?qid=20080308123211AAVM64R You can find the height by cutting the triangle in half so that one of the 2 equal sides is the hypotenuse then use Pythagorean's formula:

OpenStudy (perl):